projectile motion

A placekicker must kick a football from a point 36.0 m from the goal. Half of the crowd hopes the ball will clear the crossbar, which is 3.05 m high. When kicked,the ball leaves the ground with a speed of 20.0 m/s at an angle of 53.0 degrees to the horizontal. next stepA) By how much does the ball clear or fall short of clearing thecrossbar?

next step

B) Does the ball approach the crossbar while still rising or while falling?
Answer #1
V =20m/s
Angle is 53deg with horizontal

V(horizontal) = V(h) = 20cos53 = 12m/s
V(vertical) = V(v) = 20sin53 = 16m/s


Time taken by ball to reach 36m
T =36/12 = 3sec
Let upward direction be positive
S=uT+.5gT^2
Now the height of ball at T =3 sec =H
H = 16*3 -.5*10*3^2 = 48- 45 =3m

a)The ball fall short by 3.05-3.0 =.05 m

B) to answer this part we need to find the sign of velocity at T=3
V= u+at
=>V =16 - 10*3 = -14m/s
The ball approach the crossbar while falling since the vertical velocity becomes -ve
answered by: sanae